OpenStudy (anonymous):

integrate (x(36x^2+35)^(14)dx) if x>0

OpenStudy (anonymous):

let u = (36x^2 + 35) --> du = 72x dx do you know how to substitute them?

OpenStudy (anonymous):

yes 36/7du=14dx

OpenStudy (anonymous):

you want to substitute them to turn the integral f(x)dx to the form udu

OpenStudy (anonymous):

x dx = du/72

OpenStudy (anonymous):

ni got as my answer (36x^2+5)^15/1080

OpenStudy (anonymous):

but its saying its wrong --- i did it on my calculator

OpenStudy (anonymous):

seems right [+ C]

OpenStudy (anonymous):

yeah i also addded plus c

OpenStudy (anonymous):

it says this when i put in the answer " Can't solve for adaptive parameters"

OpenStudy (anonymous):

did you write + 35 or + 5 like you wrote just now?

OpenStudy (anonymous):

+35

OpenStudy (anonymous):

maybe the answer is infinity, having to do with the x > 0 part, but that would be weird. the integration is right though